OpenStudy (anonymous):

how to find the degree of 7% in a circle?

OpenStudy (anonymous):

plz help!!!

OpenStudy (amistre64):

what is the total degree of a circle? what is 7% of the total degree of a circle?

OpenStudy (amistre64):

spose the total degree is equal to some number, d 1% of d is equal to 1/100 of d; d/100 you want 7, 1%s sooo 7% = 7 * 1% = 7 * d/100

OpenStudy (anonymous):

360*.07=25.2